# Scheduling problem of aircrafts on a same runway and dual runways

### v3

@@ IV Takeoff scheduling on a single runway
Consider a takeoff aircraft sequence $\Phi_{a}=\langle\phi_{1},Tcf_{s_{2}},Tcf_{s_{3}},\phi_{2}\rangle$, where $\phi_{1}=\langle Tcf_{1},Tcf_{2},\cdots,Tcf_{s_{1}}\rangle$, $s_{2}=s_{1}+1$, $s_{3}=s_{1}+2$, $\phi_{2}=\langle Tcf_{s_{3}},Tcf_{s_{3}+1},\cdots,Tcf_{n}\rangle$, and $s_{1},s_{2},s_{3}$ are three positive integers. Suppose that Assumptions 1-5 hold for each landing aircraft sequence.
@@
(2) Move $Tcf_{s_{3}}$ to be between $Tcf_{h_{2}}$ and $Tcf_{h_{2}+1}$ and convert $\phi_{1}$ into a new sequence $\phi_{4}$, where $Tcf_{h_{2}},Tcf_{h_{2}+1}\in\phi_{1}$. Let $\Phi_{c}=\langle\phi_{4},Tcf_{s_{3}},\phi_{2}\rangle$. It follows that $F(\Phi_{a})-F(\Phi_{c})=\Gamma(\langle cl_{s_{2}},cl_{s_{3}},cl_{s_{3}+1},cl_{h_{2}},cl_{h_{2}+1}\rangle)=D_{cl_{s_{2}}cl_{s_{3}}}+D_{cl_{s_{3}}cl_{s_{3}+1}}+D_{cl_{h_{2}}cl_{h_{2}+1}}-D_{cl_{s_{2}}cl_{s_{3}+1}}-D_{cl_{h_{2}}cl_{s_{3}}}-D_{cl_{s_{3}}cl_{h_{2}+1}}$.
@@
Under the condition in Lemma 11, the following statements hold.
@@
(3) If $(cl_{n_{1}},cl_{n_{1}+1},j)=(\eta-1,\eta,\eta)$ for some $j\in\{1,2,\cdots,n_{1}-1\}$, then $F(\Phi_{a})-F(\Phi_{b})=0$.
@@
(3) If $h_{1}=2$, $cl_{n-1}>2$ and $\langle cl_{s_{1}},cl_{s_{1}+1}\rangle\in E_{1}$, $F(\Phi_{a})-F(\Phi_{b})=-T_{0}/3$. If $h_{1}=2$, $\langle cl_{s_{1}},cl_{s_{1}+1}\rangle\notin E_{1}$, $F(\Phi_{a})-F(\Phi_{b})<0$.
@@ V Scheduling problem of landing and takeoff aircraft on a same runway
In this section, the scheduling problem of mixed takeoff and landing aircraft on a same way for aircraft is discussed. Let $D_{T}$ denote the minimum separation time between a landing aircraft and a leading takeoff aircraft. Let $T_{D}$ denote the minimum separation time between a takeoff aircraft and a leading landing aircraft. As is defined previously, $Y_{ij}$ is used to be unified to denote the minimum separation time between any given two aircraft $Tcf_{i}$ and $Tcf_{j}$. Specifically, when a landing aircraft $Tcf_{j}$ and a leading takeoff aircraft $Tcf_{i}$ are considered, $Y_{ij}=D_{T}$; when a takeoff aircraft $Tcf_{j}$ and a leading landing aircraft $Tcf_{i}$ are considered, $Y_{ij}=T_{D}$; when two consecutive landing aircraft $Tcf_{i}$ and $Tcf_{j}$ are considered, $Y_{ij}=T_{cl_{i}cl_{j}}$; and when two consecutive takeoff aircraft $Tcf_{i}$ and $Tcf_{j}$ are considered, $Y_{ij}=D_{cl_{i}cl_{j}}$.
@@
(2) Suppose that aircraft $Tcf_{i}$ and $Tcf_{j}$ are both landing (takeoff) aircraft. If $j=i+1$ and aircraft $Tcf_{j}$ is relevant to aircraft $Tcf_{k}$, $k=i$.
@@
From Theorem 7, the optimal value of the objective function $F(\cdot)$ can be calculated along the path from from the aircraft $Tcf_{n}$ to the aircraft $Tcf_{1}$.
@@
Suppose that $Tf_{j}=[t_{0},+\infty]$ for all $j$ and Assumptions 1-4, 6-9 and 11-14 hold.
@@
(2) If the aircraft $Tcf_{b_{0}+j_{0}}$ and $Tcf_{b_{0}+j_{0}+1}$ are both takeoff or landing aircraft, $S_{(b_{0}+k)(b_{0}+k+1)}(\phi_{0})=S_{(b_{0}+k)(b_{0}+k+1)}(\phi_{1})$, for all $k=j_{0},j_{0}+1,\cdots,m-1$.

### v5

@@ IV Takeoff scheduling on a single runway
Proof: By simple calculations, this lemma can be proved and hence its proof is omitted.
@@
Consider a takeoff aircraft sequence $\Phi_{a}=\langle\phi_{1},Tcf_{s_{2}},Tcf_{s_{3}},\phi_{2}\rangle$, where $\phi_{1}=\langle Tcf_{1},Tcf_{2},\cdots,Tcf_{s_{1}}\rangle$, $s_{2}=s_{1}+1$, $s_{3}=s_{1}+2$, $\phi_{2}=\langle Tcf_{s_{3}},Tcf_{s_{3}+1},\cdots,Tcf_{n}\rangle$, and $s_{1},s_{2},s_{3}$ are three positive integers. Suppose that Assumptions 6-10 hold for each landing aircraft sequence.
@@
(2) Move $Tcf_{s_{3}}$ to be between $Tcf_{h_{2}}$ and $Tcf_{h_{2}+1}$ and convert $\phi_{1}$ into a new sequence $\phi_{4}$, where $Tcf_{h_{2}},Tcf_{h_{2}+1}\in\phi_{1}$. Let $\Phi_{c}=\langle\phi_{4},Tcf_{s_{2}},\phi_{2}\rangle$. It follows that $F(\Phi_{a})-F(\Phi_{c})=\Gamma(\langle cl_{s_{2}},cl_{s_{3}},cl_{s_{3}+1},cl_{h_{2}},cl_{h_{2}+1}\rangle)=D_{cl_{s_{2}}cl_{s_{3}}}+D_{cl_{s_{3}}cl_{s_{3}+1}}+D_{cl_{h_{2}}cl_{h_{2}+1}}-D_{cl_{s_{2}}cl_{s_{3}+1}}-D_{cl_{h_{2}}cl_{s_{3}}}-D_{cl_{s_{3}}cl_{h_{2}+1}}$.
@@
Under the conditions in Lemma 11, the following statements hold.
@@
(3) Suppose that $(cl_{n_{1}},cl_{n_{1}+1},j)=(\eta-1,\eta,\eta)$ for some $j\in\{1,2,\cdots,n_{1}-1\}$. Then $F(\Phi_{a})-F(\Phi_{b})=0$.
@@
Proof: (1)-(4) The results are obvious and hence the proofs are omitted.
(5) From Assumptions 6 and 7, $D_{11}=D_{22}=D_{21}$, $D_{33}=4T_{0}/3$ and $D_{\eta\eta}=4T_{0}/3$.
If $cl_{n_{1}}=1$, from Assumption 9, by simple calculations for the cases when $cl_{n_{1}+1}=2$, when $2<cl_{n_{1}+1}<\eta$ and when $cl_{n_{1}+1}=\eta$, it follows that $f_{1}(\Phi_{a})+f_{2}(\Phi_{a})+f_{3}(\Phi_{a})+f_{\eta}(\Phi_{a})-[f_{1}(\Phi_{b})+f_{2}(\Phi_{b})+f_{3}(\Phi_{b})+f_{\eta}(\Phi_{b})]>0$. Note that $f_{i}(\Phi_{a})=f_{i}(\Phi_{a})$ for $i\neq 1,2,3,\eta$. Hence $F(\Phi_{a})-F(\Phi_{b})>0$.
If $cl_{n_{1}}=2$, from Assumption 9, by simple calculations for the cases when $2<cl_{n_{1}+1}<\eta$ and when $cl_{n_{1}+1}=\eta$, it follows that $f_{2}(\Phi_{a})+f_{3}(\Phi_{a})+f_{\eta}(\Phi_{a})-[f_{2}(\Phi_{b})+f_{3}(\Phi_{b})+f_{\eta}(\Phi_{b})]>0$ and hence $F(\Phi_{a})-F(\Phi_{b})>0$.
If $cl_{n_{1}}=3$, from Assumption 9, by simple calculations for the cases when $3<cl_{n_{1}+1}<\eta$ and when $cl_{n_{1}+1}=\eta$, it follows that $f_{3}(\Phi_{a})+f_{\eta}(\Phi_{a})-[f_{3}(\Phi_{b})+f_{\eta}(\Phi_{b})]>0$ and hence $F(\Phi_{a})-F(\Phi_{b})>0$.
If $cl_{n_{1}}<3$, from Assumptions 9, by simple calculations for the cases when $cl_{n_{1}}<cl_{n_{1}+1}<\eta$ and when $cl_{n_{1}+1}=\eta$, it follows that $f_{cl_{n_{1}}}(\Phi_{a})+f_{\eta}(\Phi_{a})-[f_{cl_{n_{1}}}(\Phi_{b})+f_{\eta}(\Phi_{b})]>0$ and hence $F(\Phi_{a})-F(\Phi_{b})>0$.
Summarizing the above analysis, this lemma is proved.
@@
(3) If $h_{1}=2$, $cl_{n-1}>2$ and $\langle cl_{s_{1}},cl_{s_{1}+1}\rangle\in E_{1}$, $F(\Phi_{a})-F(\Phi_{b})=-T_{0}/3$. If $h_{1}=2$ and $\langle cl_{s_{1}},cl_{s_{1}+1}\rangle\notin E_{1}$, $F(\Phi_{a})-F(\Phi_{b})<0$.
@@
Proof: (1) Note that $h_{1}<cl_{s_{1}}<cl_{s_{1}+1}$. From Lemma 8, $D_{h_{1}cl_{s_{1}+1}}-D_{cl_{s_{1}}cl_{s_{1}+1}}>0.5T_{0}$. By simple calculations for the cases when $D_{cl_{s_{1}}h_{1}}=T_{0}$ and when $D_{cl_{s_{1}}h_{1}}=1.5T_{0}$, it follows that $F(\Phi_{a})-F(\Phi_{b})<0$.
(2) When $h_{1}=2$ and $\langle cl_{s_{1}},cl_{s_{1}+1}\rangle\in E_{1}$, from Lemma 8, $D_{2cl_{s_{1}+1}}-D_{cl_{s_{1}}cl_{s_{1}+1}}=T_{0}/3$. Note that $cl_{n-1}\leq 2$. Then there are at least two aircraft of class no larger than $2$ in $\phi_{2}$. By simple calculations, $F(\Phi_{a})-F(\Phi_{b})=0$.
(3) When $h_{1}=2$ and $\langle cl_{s_{1}},cl_{s_{1}+1}\rangle\in E_{1}$, from Lemma 8, $D_{2cl_{s_{1}+1}}-D_{cl_{s_{1}}cl_{s_{1}+1}}=T_{0}/3$. Note that $cl_{n-1}>2$. Then there is at most one aircraft of class no larger than $2$ in $\phi_{2}$. By simple calculations, $F(\Phi_{a})-F(\Phi_{b})=\Gamma(cl_{i},h_{1},cl_{i+1},cl_{s_{1}},cl_{s_{1}+1})=-T_{0}/3$. When $h_{1}=2$ and $\langle cl_{s_{1}},cl_{s_{1}+1}\rangle\in E_{1}$, from Lemma 8, $D_{2cl_{s_{1}+1}}-D_{cl_{s_{1}}cl_{s_{1}+1}}>0.5T_{0}$. By simple calculations, $F(\Phi_{a})-F(\Phi_{b})<0$.
The statements (4)-(9) can be proved by using similar approaches in (1)-(3).
@@ V Scheduling problem of landing and takeoff aircraft on a same runway
In this section, the scheduling problem of mixed takeoff and landing aircraft on a same way for aircraft is discussed. Let $D_{T}$ denote the minimum separation time between a landing aircraft and a leading takeoff aircraft. Let $T_{D}$ denote the minimum separation time between a takeoff aircraft and a leading landing aircraft. As defined previously, $Y_{ij}$ is used to unifiedly denote the minimum separation time between any given two aircraft $Tcf_{i}$ and $Tcf_{j}$. Specifically, when a landing aircraft $Tcf_{j}$ and a leading takeoff aircraft $Tcf_{i}$ are considered, $Y_{ij}=D_{T}$; when a takeoff aircraft $Tcf_{j}$ and a leading landing aircraft $Tcf_{i}$ are considered, $Y_{ij}=T_{D}$; when two consecutive landing aircraft $Tcf_{i}$ and $Tcf_{j}$ are considered, $Y_{ij}=T_{cl_{i}cl_{j}}$; and when two consecutive takeoff aircraft $Tcf_{i}$ and $Tcf_{j}$ are considered, $Y_{ij}=D_{cl_{i}cl_{j}}$.
@@
(2) Suppose that aircraft $Tcf_{i}$ and $Tcf_{j}$ are both landing (takeoff) aircraft. If $j=i+1$ and aircraft $Tcf_{j}$ is relevant to aircraft $Tcf_{k}$, then $k=i$.
@@
Proof: (1) Since $Y_{ij}=D_{T}<1.5T_{0}$ from Assumption 11, if $j\geq i+2$, then $S_{ij}\geq 2T_{0}>Y_{ij}$ and aircraft $Tcf_{j}$ is not relevant to aircraft $Tcf_{i}$. If aircraft $Tcf_{j}$ is relevant to aircraft $Tcf_{i}$, then $j=i+1$.
(2) Note that $D_{T}<1.5T_{0}$ and $T_{D}<1.5T_{0}$ from Assumption 11. If $k\neq i$, it follows that $k<i$, and $Tcf_{k}$ is a landing (takeoff) aircraft. From Lemma 1, $S_{kj}\geq S_{ij}+S_{ki}\geq Y_{i{j}}+Y_{k{i}}>Y_{k{j}}$, implying that aircraft $Tcf_{j}$ is not relevant to aircraft $Tcf_{k}$. Therefore, $k=i$.
(3)-(4) These two is obvious and hence their proofs are omitted.
(5) It is clear that the aircraft $Tcf_{i}$ and $Tcf_{j}$ are both landing (or takeoff) aircraft, and $Y_{ij}=3T_{0}$. This implies that $Y_{(j-1)j}=T_{0}$ and $Y_{(j-2)j}=2T_{0}$. From (3), the aircraft $Tcf_{j-1}$ and $Tcf_{j-2}$ are both takeoff (or landing) aircraft. Thus, the aircraft $Tcf_{j}$ is also relevant to aircraft $Tcf_{j-1}$ and not relevant to aircraft $Tcf_{j-2}$.
@@
Proof: (1) Sufficiency is obvious and we only discuss the necessity of this theorem. Suppose that there is no path from the aircraft $Tcf_{n}$ to the aircraft $Tcf_{1}$, when the optimization problem (6) is solved. Construct an aircraft set $\Theta_{1}$ such that there is a path from the aircraft $Tcf_{n}$ to the aircraft $Tcf_{i}$ for any $Tcf_{i}\in\Theta_{1}$ and there is no path from the aircraft $Tcf_{n}$ to the aircraft $Tcf_{i}$ for any $Tcf_{i}\in\{Tcf_{1},Tcf_{2},\cdots,Tcf_{n}\}-\Theta_{1}$. Shift the takeoff or landing times of all the aircraft in $\Theta_{1}$ as a whole to an earlier time instant with their relative landing or takeoff times unchanged while keeping necessary minimum separation times with the aircraft in $\{Tcf_{1},Tcf_{2},\cdots,Tcf_{n}\}-\Theta_{1}$. Such a operation can be realized because the aircraft in $\Theta_{1}$ are not relevant to any aircraft in $\{Tcf_{1},Tcf_{2},\cdots,Tcf_{n}\}-\Theta_{1}$. After this operation, the value of the objective function $F(\phi,Sr(\phi))$ becomes smaller, which yields a contraction. Therefore, there is a path from the aircraft $Tcf_{n}$ to the aircraft $Tcf_{1}$.
(2) From Theorem 6 and the definition of path, if the aircraft $Tcf_{2}$ is relevant to the aircraft $Tcf_{1}$ and the aircraft $Tcf_{i+2}$ is relevant to the aircraft $Tcf_{i}$ or $Tcf_{i+1}$, $i=1,2,\cdots,n-2$, there is a path from the aircraft $Tcf_{n}$ to the aircraft $Tcf_{1}$.
@@
From Theorem 7, the optimal value of the objective function $F(\cdot)$ can be calculated along the path from the aircraft $Tcf_{n}$ to the aircraft $Tcf_{1}$.
@@
Suppose that $Tf_{j}=[t_{0},+\infty]$ for all $j$, and Assumptions 1-4, 6-9 and 11-14 hold.
@@
Proof: If $\max\{S_{(i+1)(i+2)}(\phi_{0}^{1}),S_{(i+1)(i+2)}(\phi_{0}^{2})\}=Y_{{(i+1)}{(i+2)}}$, then in $\phi_{0}^{1}$ and $\phi_{0}^{2}$, the aircraft $Tcf_{i+2}$ is relevant to the aircraft $Tcf_{i+1}$, and $S_{(i+1)(i+2)}(\phi_{0}^{1})=S_{(i+1)(i+2)}(\phi_{0}^{2})$. Suppose that $\max\{S_{(i+1)(i+2)}(\phi_{0}^{1}),S_{(i+1)(i+2)}(\phi_{0}^{2})\}$ $>Y_{{(i+1)}{(i+2)}}$, and $S_{(i+1)(i+2)}(\phi_{0}^{1})\geq S_{(i+1)(i+2)}(\phi_{0}^{2})$. Clearly, $S_{(i+1)(i+2)}(\phi_{0}^{1})>Y_{({i+1})({i+2})}$. In $\phi_{0}^{1}$ and $\phi_{0}^{2}$, the aircraft $Tcf_{i+2}$ is relevant to the aircraft $Tcf_{i}$ or the aircraft $Tcf_{i+1}$. It follows that $S_{i(i+2)}(\phi_{0}^{1})=Y_{{i}({i+2})}$. On the other hand, note that $S_{i(i+2)}(\phi_{0}^{1})=Y_{{i}({i+2})}>S_{(i+1)(i+2)}(\phi_{0}^{2})+S_{i(i+1)}(\phi_{0}^{2})$, which contradicts with the definition of relevance. Therefore, $\max\{S_{(i+1)(i+2)}(\phi_{0}^{1}),S_{(i+1)(i+2)}(\phi_{0}^{2})\}=Y_{({i+1})({i+2})}$, or $S_{(i+1)(i+2)}(\phi_{0}^{1})<S_{(i+1)(i+2)}(\phi_{0}^{2})$.
When $S_{(i+1)(i+2)}(\phi_{0}^{1})<S_{(i+1)(i+2)}(\phi_{0}^{2})$, if $S_{(i+1)(i+2)}(\phi_{0}^{2})=Y_{(i+1)(i+2)}$, then $Y_{(i+1)(i+2)}>S_{(i+1)(i+2)}(\phi_{0}^{1})$, which also contradicts with the definition of relevance. Note that in $\phi_{0}^{1}$ and $\phi_{0}^{2}$, the aircraft $Tcf_{i+2}$ is relevant to the aircraft $Tcf_{i}$ or the aircraft $Tcf_{i+1}$. It follows that $S_{i(i+2)}(\phi_{0}^{2})=Y_{i(i+2)}$.
@@
(2) If the aircraft $Tcf_{b_{0}+j_{0}}$ and $Tcf_{b_{0}+j_{0}+1}$ are both takeoff or landing aircraft, $S_{(b_{0}+k)(b_{0}+k+1)}(\phi_{0})=S_{(b_{0}+k)(b_{0}+k+1)}(\phi_{1})$ for all $k=j_{0},j_{0}+1,\cdots,m-1$.
@@
Proof: (1) From Assumption 14, the statement (1) holds.
(2) If the aircraft $Tcf_{b_{0}}$ and $Tcf_{b_{0}+1}$ are both takeoff or landing aircraft, from Theorem 6, in $\phi_{0}$ and $\phi_{1}$, the aircraft $Tcf_{b_{0}+1}$ is relevant to the aircraft $Tcf_{b_{0}}$, and $S_{b_{0}(b_{0}+1)}(\phi_{0})=S_{b_{0}(b_{0}+1)}(\phi_{1})$. From the statement (1), the statement (2) holds.
(3) In view of the condition that $S_{b_{0}(b_{0}+1)}(\phi_{0})\neq S_{b_{0}(b_{0}+1)}(\phi_{1})$, without loss of generality, suppose that $S_{b_{0}(b_{0}+1)}(\phi_{0})>S_{b_{0}(b_{0}+1)}(\phi_{1})$. It follows that $S_{b_{0}(b_{0}+1)}(\phi_{0})>Y_{b_{0}(b_{0}+1)}$. That is, the aircraft $Tcf_{b_{0}+1}$ is not relevant to the aircraft $Tcf_{b_{0}}$ in $\phi_{0}$. From Assumption 14, the aircraft $Tcf_{b_{0}+1}$ is relevant to the last aircraft of $\phi_{01}$. If $S_{(b_{0}+1)(b_{0}+2)}(\phi_{0})=S_{(b_{0}+1)(b_{0}+2)}(\phi_{1})$, from the statement (1), $S_{(b_{0}+k)(b_{0}+k+1)}(\phi_{0})=S_{(b_{0}+k)(b_{0}+k+1)}(\phi_{1})$ for all $k=1,2,\cdots,m-1$. Suppose that $S_{(b_{0}+1)(b_{0}+2)}(\phi_{0})\neq S_{(b_{0}+1)(b_{0}+2)}(\phi_{1})$, from Lemma 15, $S_{(b_{0}+1)(b_{0}+2)}(\phi_{0})<S_{(b_{0}+1)(b_{0}+2)}(\phi_{1})$, and the aircraft $Tcf_{b_{0}+2}$ is relevant to $Tcf_{b_{0}}$ in $\phi_{1}$. By analogy, if $S_{(b_{0}+2)(b_{0}+3)}(\phi_{0})=S_{(b_{0}+2)(b_{0}+3)}(\phi_{1})$, then $S_{(b_{0}+k)(b_{0}+k+1)}(\phi_{0})=S_{(b_{0}+k)(b_{0}+k+1)}(\phi_{1})$ for all $k=2,3,\cdots,m-1$. When $S_{(b_{0}+2)(b_{0}+3)}(\phi_{0})\neq S_{(b_{0}+2)(b_{0}+3)}(\phi_{1})$, from Lemma 15, $S_{(b_{0}+2)(b_{0}+3)}(\phi_{0})>S_{(b_{0}+2)(b_{0}+3)}(\phi_{1})$, and the aircraft $Tcf_{b_{0}+3}$ is relevant to $Tcf_{b_{0}+1}$ in $\phi_{0}$.
Note that the aircraft $Tcf_{b_{0}+1}$ is relevant to the last aircraft of $\phi_{01}$ in $\phi_{0}$ and the aircraft $Tcf_{b_{0}+3}$ is relevant to the aircraft $Tcf_{b_{0}+1}$ in $\phi_{0}$. From Assumption 12, the aircraft $Tcf_{b_{0}+5}$ is relevant to the aircraft $Tcf_{b_{0}+4}$ in $\phi_{0}$ and $Y_{(b_{0}+3)(b_{0}+5)}<T_{D}+D_{T}$. Since $Tcf_{b_{0}+4}$ is a takeoff (landing) aircraft and $Tcf_{b_{0}+3},Tcf_{b_{0}+5}$ are landing (takeoff) aircraft, $S_{(b_{0}+3)(b_{0}+5)}(\phi_{1})\geq T_{D}+D_{T}>Y_{(b_{0}+3)(b_{0}+5)}$. Therefore, the aircraft $Tcf_{b_{0}+5}$ is also relevant to the aircraft $Tcf_{b_{0}+4}$ in $\phi_{1}$. From statement (1), $S_{(b_{0}+k)(b_{0}+k+1)}(\phi_{0})=S_{(b_{0}+k)(b_{0}+k+1)}(\phi_{1})$ for all $k=4,5,\cdots,m-1$.